OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? 2x-5y=-5 x+2y=11 A. (3, 4) B. (5, 3) C. (7, 2) D. (15, 7)

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Have you tried it? :P

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i think b or d

OpenStudy (luigi0210):

Plug them in and check: For (5, 3): 2(5)-5(3)=-5 (5)+2(3)=11 If they turn out true, then they are your answer

OpenStudy (owllover101):

i think b

OpenStudy (anonymous):

thats what i put

OpenStudy (luigi0210):

@owllover101 Don't give away answers >_>

OpenStudy (anonymous):

haha sooo i was right ?

OpenStudy (luigi0210):

if you did it right it should be